# Experimental Ground-State Energy of a 125-Site Flat Kagome Antiferromagnet via Hamiltonian Engineering on Quantum Computer

### v2

@@ Abstract
We present an instance of utility-grade quantum computation by calculating the ground-state energy of a 125-site flat Kagome lattice under antiferromagnetic Heisenberg model (KAFH), with IBM’s Heron r1 and Heron r2 quantum processors. For spin-1/2 KAFH, our best per-site ground-state energy estimate arrives at $-0.417\,J$, and after applying open-boundary corrections, it closely approaches the established thermodynamic value of $-0.438\,J$. To achieve this, we propose a hybrid approach that splits the variational quantum eigensolver (VQE) into local (classical) and global (quantum) components for efficient hardware utilization. We further introduce a Hamiltonian engineering strategy that increases coupling on defect triangles to mimic loop-flip dynamics, allowing us to simplify the ansatz while retaining physical accuracy. Using a single-repetition, hardware-efficient ansatz, we entangle up to 103 qubits with high fidelity to determine the Hamiltonian lowest eigenvalue. This work demonstrates the scalability of VQE for frustrated 2D systems and lays the foundation for the future studies using deeper ansatz circuits and larger lattices.
@@ I Introduction
Recent advances in quantum hardware have brought general-purpose quantum processors to the threshold of solving problems that challenge classical computation, particularly in quantum chemistry and strongly correlated materials [1, 2]. Among the most promising tools is the Variational Quantum Eigensolver (VQE) [15], a hybrid quantum-classical algorithm well-suited to current noisy intermediate-scale quantum (NISQ) devices [33], especially those following the heavy-hex architecture, such as IBM’s Falcon and Hummingbird processors [3, 4]. With public access to devices supporting over 100 qubits, VQE is being extended to simulate increasingly large systems. However, the translation of reduced noise and improved coherence times into utility-grade quantum computation remains an open question, especially in challenging domains like quantum magnetism, where ground-state preparation and energy estimation are complicated by frustration and topological order [5, 6]. The central aim of this study lies in the multi-prong approach for the utility grade quantum computing, a crucial milestone on the path to quantum advantage [37].
@@
Figure 1: A proof-of-concept demonstration of Hamiltonian engineering for lowering the ground-state energy. (a) A 6-site subregion of the Kagome lattice targeted for ground-state energy calculation. (b) The corresponding 6-qubit VQE ansatz circuit containing chain of CNOT gates sandwiched between parametric $R_{y}(\theta)$ gates. (c) Ground-state energy (GSE) and bond energy map $\textnormal{E}_{\texttt{XY}}$ without Hamiltonian engineering, showing localized static dimers. (d) Applying Hamiltonian engineering with enhanced coupling $J^{\prime}=2$ lowers the GSE. The bond energy map shows that the static dimers (q2, q3) and (q4, q5) evolve into superposition of dimers, entangling all four qubits, thus simulating the spin-liquid state.
@@
In this work, we showcase a scalable VQE framework to estimate the ground-state energy of the 125-site spin-1/2 Kagome Antiferromagnetic Heisenberg model (KAFH) on a fully flat 2D lattice shown in Figure-2. Note that throughout this manuscript, we assume $J=1$ and, unless otherwise specified, report four times (4x) the ground-state energy for convenience in calculations and data presentation. This means that singlet energy is taken -3 instead of -3/4. Unlike prior studies constrained to quasi-1D geometries (e.g., cylinders or strips) for computational tractability [7], our method preserves the full 2D translational and topological symmetry of the lattice. We introduce a Hamiltonian engineering strategy: a valence bond crystal (VBC) dimer configuration is used as a physically motivated starting point, and local exchange couplings are modified (specifically, increasing one bond of each defect triangle to $J^{\prime}\approx 2$, while keeping all others at $J=1$) to simulate the energetic effects of loop-flips and drive dimer resonance, as highlighted in Figure-1. This adjustment compensates the relatively shallow, hardware-efficient ansatz [3] for their lower expressiveness, particularly our single-repetition real-amplitude circuit supporting only 1D nearest-neighbor entanglement, and helps induce quantum fluctuations in the vicinity of defect triangles, enabling a superposition of dimer covers reminiscent of the resonating valence bond (RVB) spin-liquid picture [8, 9].
@@
To implement this approach efficiently, we partition our VQE algorithm into local and global phases. The local VQE is executed on classical hardware, decomposing the 103-qubit ansatz (Figure-6(a)) into overlapping 15–19 qubit subcircuits that each span defect triangles and nearby corner-sharing structures. Each subcircuit is independently optimized, with $J^{\prime}$ adjusted to match known ground-state energies from exact diagonalization (feasible at these small scales). The subcircuits are then stitched into a complete ansatz with additional parametric single-qubit gates at the interfaces, see Figure-3. The global VQE phase runs on the quantum processor to optimize only these bridging gates, potentially capturing longer-range entanglement. Final energy estimates are refined using the Operator Decoherence Renormalization (ODR) [30] noise-mitigation protocol [10]. This work pushes the frontier of practical quantum simulation by: (1) offloading expensive local optimization to classical hardware (2) enabling high-fidelity energy estimation through localized Hamiltonian engineering (3) demonstrating progress toward utility-scale quantum computation (4) preserving the geometric and topological integrity of the Kagome lattice during simulation.
@@
To our knowledge, this is the first application of local Hamiltonian calibration—via bond-specific $J$ tuning—to enhance VQE performance on large 2D Kagome lattices. In contrast to previous VQE or hybrid tensor-network approaches [21, 22, 24], which often rely on quasi-1D or complete-graph hardware, our method preserves full 2D topology and directly engineers local fluctuations crucial for RVB and topological order. We believe this technique opens a new avenue for scalable quantum simulation of frustrated magnetism and spin-liquid physics.
The manuscript organization discusses prior work in Section-II, followed by the details of hamiltonian engineering in Section-III. The proposed VQE algorithm is described in Section-IV while concomitant experimental results are provided in Section-V. The Section-VI validates our experimentally obtained gs_energy estimate and Section-VII concludes the article with possible directions for the future studies.
@@ II Prior Work
The quest to accurately determine the ground state properties of strongly correlated quantum many-body systems remains a central challenge in condensed matter physics. Among these, geometrically frustrated magnets, such as the Kagome Antiferromagnet (KAFM), are of particular interest due to their potential to host exotic quantum phases, notably the quantum spin liquid (QSL) [8, 11]. The spin-1/2 Heisenberg KAFH is theoretically predicted to exhibit a gapped topological $\mathbb{Z}_{2}$ spin liquid ground state, characterized by robust long-range entanglement and fractionalized excitations [12, 13, 14]. Classical numerical techniques like Density Matrix Renormalization Group (DMRG) [35] and exact diagonalization (ED) have provided deep insights for smaller systems, but their scalability limitations restrict exploration of the thermodynamic regime.
Figure 2: (Color online) The 125-site flat Kagome(color online) The 125-site flat Kagome lattice showing static dimer coverings (thick black line-segments) and defect triangles (in red), driven by the our qubit-lattic site map. The numerical labels on the sites correspond to qubit indices in the 103-qubit ansatz circuit depicted in Figure 6(a). These qubits were included in the quantum computation of the ground-state energy gs_energy. Alphabetically labeled sites were excluded from the quantum circuit but were included in the final per-site ground-state energy calculation. The dotted region was omitted entirely and not mapped to the quantum problem in this study.
VQE has emerged as a scalable alternative [3, 15, 16], enabling ground-state estimation on NISQ devices. Early work applied VQE to 1D Heisenberg models [17], followed by studies of frustrated systems [18, 19]. Jattana et al. [20] explored the 2D Heisenberg model on square and triangular lattices using VQE. More recent studies investigated frustrated lattices like the Kagome: Kattemölle and van Wezel [21] proposed SU(2)-symmetric ansätze; Bosse and Montanaro [22] demonstrated $<1\%$ energy error on small Kagome patches; Guo and Qin [23] examined J1-J2 phase transitions; and Javanmard et al. [24] implemented a hybrid MPS-VQE approach for noise-resilient circuits. Experimental implementations have also progressed: in Ref [25], competition, winners reported sub-1% energy error on IBM hardware, and Grossardt et al. [26] showed contextual VQE can achieve 0.01% accuracy for small Kagome clusters. Yet, full 2D Kagome connectivity on large quantum devices remains unexplored—until now.
@@ III Hamiltonian Engineering on the Defect Triangles
Quantum spin liquids, emergent phases characterized by long-range entanglement and a highly resonant superposition of valence bond (dimer) configurations on frustrated lattices, have attracted intense theoretical and experimental scrutiny over the past decade [11, 5, 27]. The Kagome antiferromagnetic Heisenberg (KAFH) model is a prototypical platform for realizing such physics, with compelling evidence, via DMRG and quantum dimer model methods, favoring a gapped $\mathbb{Z}_{2}$ spin liquid ground state [7, 13, 12]. However, nearly all high-precision simulations employ quasi-1D geometries (cylinders or strips), which necessarily break the lattice’s 2D translational and topological symmetries. Here, we propose a scalable quantum computational strategy based on a 100-site square Kagome lattice, preserving the genuine 2D lattice topology.
@@ III.1 Calibration the Hamiltonian
Our approach begins with a static valence bond covering—a natural reference state possessing one defect triangle per four total triangles. To remediate the systematic under-representation of dimers on defect sites in the static picture, we tune one Heisenberg coupling on each defect triangle from $J=1$ to $J^{\prime}=2$. This calibration aligns the energy gain from creating a new dimer with the energy penalty of removing two existing dimers, as one expects from local energy conservation arguments in RVB or dimer models. It injects targeted quantum fluctuations—loop resonances—around defect triangles, encouraging the system toward spin-liquid-like superposition as showcased in Figure-1.
This localized increase in coupling aims to strongly induce dimerization within these specific triangles, making it energetically favorable for our ansatz to form a singlet. This targeted modification results in an average interaction strength of $J_{\text{avg}}=\frac{1+1+2}{3}=\frac{4}{3}$ on these defect triangles. By energetically penalizing non-dimerized configurations in these frustrated regions, we compel the VQE to explore quantum states closer to the Kagome ground state, aiming to match its energy. Our 125-site Kagome patch, with open boundaries in both $X$ and $Y$ directions, introduces additional finite-size effects and results in a higher (less negative) ground-state energy compared to the thermodynamic limit. However, the proposed Hamiltonian engineering focuses on enabling our ansatz to efficiently discover a low-energy configuration within this finite system, thereby approximating the physics of the larger Kagome lattice.
More importantly, the calibrated Hamiltonian approach leverages a hardware-efficient Variational Quantum Eigensolver (VQE) framework. Our ansatz is the Qiskit [31] Real Amplitude circuit. Figure-6(a) shows that it is 1-D chain of CNOT entanglers interleaved with parameterized single-qubit rotations, designed for compatibility with IBM’s heavy-hex superconducting architecture. This circuit can entangle nearest-neighbor qubits and is sufficiently expressive to capture local resonance physics, while remaining shallow and hardware-friendly.
Table 1: Effectiveness of Hamiltonian engineering with $J^{\prime}\approx 2$ in matching the estimated ground-state energy (gs_energy) to the exact value for progressively larger subregions of the Kagome lattice, as shown in Figure 7.
No. of Sites Subregions of Kagome Lattice of Figure-7 No. of Defect Triangles Exact gs_energy (units of J) VQE computed gs_energy (units of J) J’ (units of J)
@@ III.2 Scalability of Hamiltonian Calibration
Table 1 summarizes example Hamiltonian engineering for various Kagome lattice subregions containing 6 to 23 sites. For each subregion size $N$, classical VQE simulations with $J^{\prime}\approx 2$ yield ground-state energies in close agreement with exact diagonalization. The corresponding geometries are depicted in Figure 7, each chosen to ensure that at least one defect triangle is present and that the ansatz qubit mapping accommodates local entanglement. In some configurations (e.g., $N=23$), two or more defect triangles are included. In each sublattice, the specific bond on the defect triangle where the Heisenberg coupling was approximately doubled is marked as $J^{\prime}$. The data highlights that the same calibrated interaction strength $J^{\prime}$ suffices across different lattice geometries without the need for re-tuning, reinforcing the scalability of the method.
@@
III.3 Physical Interpretation of J’ = 2
The smallest loop in the Kagome lattice that allows dimer flipping is the hexagonal plaquette contained within a single unit cell. We justify setting $J^{\prime}=2$ by demonstrating that, at this value, the ansatz circuit enables a complete loop-flip and reproduces the exact ground-state energy. Figure-7(c) shows a 18-site Kagome unit cell whose gs_energy was computed from loop version of 6-qubit hardware-efficient ansatz circuit, which applies CNOT 5,0 to complete the loop of entangling gates. The initial dimer configuration places static singlets on edges (0,1), (2,3), and (4,5), leading to an overestimated energy of gs_energy $=-9$, while leaving three defect triangles. When the exchange coupling $J^{\prime}=2$ is applied on bonds (1,2), (3,4), and (5,0), the ansatz induces a loop-flip within the central hexagon and recovers the exact ground-state energy gs_energy $=-18$. This proof-of-concept validates that targeted Hamiltonian engineering on defect triangles can correct energy errors arising from suboptimal qubit mappings or limited ansatz expressiveness.
Table 2: Components of the ground-state energy (gs_energy) computed using local VQE. Each component corresponds to a subregion shown in Figure 8, with its respective defect triangle bond and its fine-tuned $J^{\prime}$.
Lattice subregion of the Figure-7 Bonds with J’ J’ (units of J) Local VQE Sub-lattice energy (units of J)

### v3

# Utility-Scale Quantum Computation of Ground-State Energy in a 100+ Site Planar Kagome Antiferromagnet via Hamiltonian Engineering

@@ Abstract
We present experimental quantum computation of the ground-state energy in a 103-site flat Kagome lattice under antiferromagnetic Heisenberg model (KAFH), with IBM’s Heron r1 and Heron r2 quantum processors. For spin-1/2 KAFH, our per-site ground-state energy estimate is $-0.417\,J$, which, under open-boundary corrections, matches the energy in thermodynamic limit. i.e. $-0.4386\,J$. To achieve this, we used a hybrid approach that splits the conventional Variational Quantum Eigensolver (VQE) into local (classical) and global (quantum) components for efficient hardware utilization. More importantly, we introduce a Hamiltonian engineering strategy that increases coupling on defect triangles to mimic loop-flip dynamics, allowing us to simplify the ansatz while retaining computational accuracy. Using a single-repetition, hardware-efficient ansatz, we entangle up to 103 qubits with high fidelity to determine the Hamiltonian lowest eigenvalue. This work demonstrates the scalability of VQE for frustrated 2D systems and lays the foundation for the future studies using deeper ansatz circuits and larger lattices on utility quantum processors.
@@ I Introduction
Recent advances in quantum hardware have brought general-purpose quantum processors to the threshold of solving problems that challenge classical computation, particularly in quantum chemistry and strongly correlated materials [1, 2]. Among the most promising tools is the Variational Quantum Eigensolver (VQE) [15], a hybrid quantum-classical algorithm well-suited to current noisy intermediate-scale quantum (NISQ) devices [33], especially those following the heavy-hex architecture, such as IBM’s Falcon and Hummingbird processors [3, 4]. With public access to devices supporting over 100 qubits, VQE is being extended to simulate increasingly large systems. However, the translation of reduced noise and improved coherence times into utility-scale quantum computation remains an open question, especially in challenging domains like quantum magnetism, where ground-state preparation and energy estimation are complicated by frustration and topological order [5, 6]. The central aim of this study lies in the multi-prong approach for the utility-scale quantum computing [38, 39, 40, 41, 42, 30], a crucial milestone on the path to quantum advantage [37].
@@
Figure 1: A proof-of-concept demonstration of Hamiltonian engineering for lowering the ground-state energy. (a) A 6-site subregion of the Kagome lattice targeted for ground-state energy calculation. (b) The corresponding 6-qubit VQE ansatz circuit containing chain of CNOT gates sandwiched between parametric $R_{y}(\theta)$ gates. (c) Ground-state energy (gs_energy) and bond energy map $\textnormal{e}_{\texttt{XY}}$ without Hamiltonian engineering, showing localized static dimers. (d) Applying Hamiltonian engineering with enhanced coupling $J^{\prime}=2$ lowers the gs_energy. The bond energy map shows that the static dimers (q2, q3) and (q4, q5) evolve into superposition of dimers, entangling all four qubits, thus simulating the spin-liquid state.
@@
In this work, we showcase a scalable VQE framework to estimate the ground-state energy (gs_energy) of the 103-site spin-1/2 Kagome Antiferromagnetic Heisenberg model (KAFH) on a fully flat 2D lattice shown in Figure-2. Note that throughout this manuscript, we assume exchange coupling strength $J=1$ and, unless otherwise specified, report four times (4x) the ground-state energy for convenience in calculations and data presentation. This means that singlet energy is taken $-3$ instead of $-3/4$. Unlike prior studies constrained to quasi-1D geometries (e.g., cylinders or strips) for computational tractability [7], our method preserves the full 2D translational and topological symmetry of the lattice. We introduce a Hamiltonian engineering strategy: a valence bond crystal (VBC) dimer configuration is used as a physically motivated starting point, and local exchange couplings are modified (specifically, increasing one bond of each defect triangle to $J^{\prime}\approx 2$, while keeping all others at $J=1$) to simulate the energetic effects of loop-flips and drive dimer resonance, as highlighted in Figure-1. This adjustment compensates the relatively shallow, hardware-efficient ansatz [3] for their lower expressiveness, particularly our single-repetition real-amplitude circuit supporting only 1D nearest-neighbor entanglement, and helps induce quantum fluctuations in the vicinity of defect triangles, enabling a superposition of dimer covers reminiscent of the resonating valence bond (RVB) spin-liquid picture [8, 9].
@@
To implement this approach efficiently, we partition our VQE algorithm into local and global phases. The local VQE is executed on classical hardware, decomposing the 103-qubit ansatz (Figure-6(a)) into overlapping 15–19 qubit subcircuits that each span defect triangles and nearby corner-sharing structures. Each subcircuit is independently optimized, with $J^{\prime}$ adjusted to match known ground-state energies from exact diagonalization (feasible at these small scales). The subcircuits are then stitched into a complete ansatz with additional parametric single-qubit gates at the interfaces, see Figure-3. The global VQE phase runs on the quantum processor to optimize only these bridging gates, potentially capturing longer-range entanglement. Final energy estimates are refined using the Operator Decoherence Renormalization (ODR) [30] noise-mitigation protocol [10]. This work pushes the frontier of practical quantum simulation by: (1) offloading expensive local optimization to classical hardware (2) enabling high-fidelity energy estimation through localized Hamiltonian engineering (3) demonstrating utility-scale quantum computation (4) preserving the geometric and topological integrity of the Kagome lattice during simulation.
@@
To our knowledge, this is the first application of local Hamiltonian calibration, via bond-specific $J$ tuning, to enhance VQE performance on large 2D Kagome lattices. In contrast to previous VQE or hybrid tensor-network approaches [21, 22, 24], which often rely on quasi-1D or complete-graph hardware, our method preserves full 2D topology and directly engineers local fluctuations crucial for RVB and topological order. We believe this technique opens a new avenue for scalable quantum simulation of frustrated magnetism and spin-liquid physics. The manuscript organization situates the prior work in Section-II, followed by the details of hamiltonian engineering in Section-III. The proposed VQE algorithm is described in Section-IV while experimental results are provided in Section-V. The Section-VI validates our experimentally obtained gs_energy estimate and Section-VII concludes the article with possible directions for the future studies. The Qiskit code and related material is available at https://github.com/ahsan-quantum/Kagome-Lattice-VQE.git
@@ II Prior Work
The quest to accurately determine the ground state properties of strongly correlated quantum many-body systems remains a central challenge in condensed matter physics. Among these, geometrically frustrated magnets, such as the Kagome Antiferromagnet, are of particular interest due to their potential to host exotic quantum phases, notably the quantum spin liquid (QSL) [8, 11]. The spin-1/2 Heisenberg KAFH is theoretically predicted to exhibit a gapped topological $\mathbb{Z}_{2}$ spin liquid ground state, characterized by robust long-range entanglement and fractionalized excitations [12, 13, 14]. Classical numerical techniques like Density Matrix Renormalization Group (DMRG) [35] and exact diagonalization (ED) have provided deep insights for smaller systems, but their scalability limitations restrict exploration of the thermodynamic regime.
Figure 2: (Color online) The 125-site planar Kagome lattice showing static dimer coverings (thick black line-segments) and defect triangles (in red), driven by the our qubit-lattic site map. The numerical labels on the sites correspond to qubit indices in the 103-qubit ansatz circuit depicted in Figure 6(a). These qubits were included in the quantum computation of the gs_energy of 103-site KAFH. Alphabetically labeled sites were excluded from the quantum circuit but were included in the final per-site ground-state energy calculation. The dotted region was omitted entirely and not mapped to the quantum problem in this study.
VQE has emerged as a scalable alternative [3, 15, 16], enabling ground-state estimation on NISQ devices. Early work applied VQE to 1D Heisenberg models [17], followed by studies of frustrated systems [18, 19]. Jattana et al. [20] explored the 2D Heisenberg model on square and triangular lattices using VQE. More recent studies investigated frustrated lattices like the Kagome: Kattemölle and van Wezel [21] proposed SU(2)-symmetric ansätze; Bosse and Montanaro [22] demonstrated $<1\%$ energy error on small Kagome patches; Guo and Qin [23] examined J1-J2 phase transitions; and Javanmard et al. [24] implemented a hybrid MPS-VQE approach for noise-resilient circuits. Experimental implementations have also progressed: in Ref [25], competition, winners reported sub-1% energy error on IBM hardware, and Grossardt et al. [26] showed contextual VQE can achieve 0.01% accuracy for small Kagome clusters. Yet, full 2D Kagome connectivity on large quantum devices has remained largely unexplored.
@@ III Hamiltonian Engineering on the Defect Triangles
Quantum spin liquids, emergent phases characterized by long-range entanglement and a highly resonant superposition of valence bond (dimer) configurations on frustrated lattices, have attracted intense theoretical and experimental scrutiny over the past decade [11, 5, 27]. The KAFH model is a prototypical platform for realizing such physics, with compelling evidence, via DMRG and quantum dimer model methods, favoring a gapped $\mathbb{Z}_{2}$ spin liquid ground state [7, 13, 12]. However, nearly all high-precision simulations employ quasi-1D geometries (cylinders or strips), which necessarily break the lattice’s 2D translational and topological symmetries. Here, we propose a scalable quantum computational strategy based on a 125-site planar Kagome lattice, preserving the intrinsic 2D lattice topology.
@@ III.1 Calibration the Hamiltonian
Let a $\mathcal{K}=\{(\alpha,\beta)\}$ denote set of all pairs ($\alpha,\beta$) representing neighboring sites of the Kagome lattice shown in Figure-2. Accordingly the KAFH Hamiltonian $H$ becomes
$H=J\sum_{(\alpha,\beta)\in\mathcal{K}}X_{\alpha}X_{\beta}+Y_{\alpha}Y_{\beta}+Z_{\alpha}Z_{\beta}$
and $H_{pert}$ is the engineered hamiltonian which enhances the exchange coupling $J\longrightarrow J^{\prime}$ on some $\mathcal{S}\in\mathcal{K}$ to be defined on the defect triangles. Then,
$H_{pert}=\;$ $J^{\prime}_{(\alpha,\beta)}\sum_{(\alpha,\beta)\in\mathcal{S}}\left(X_{\alpha}X_{\beta}+Y_{\alpha}Y_{\beta}+Z_{\alpha}Z_{\beta}\right)$
$+J\sum_{(\alpha,\beta)\in\mathcal{K}\setminus\mathcal{S}}\left(X_{\alpha}X_{\beta}+Y_{\alpha}Y_{\beta}+Z_{\alpha}Z_{\beta}\right)$ (1)
Our approach begins with a static valence bond covering, a natural reference state creating one defect triangles as shown in Figure-2. To remediate the systematic under-representation of dimers on defect sites in the static picture, we tune one Heisenberg coupling on each defect triangle from $J=1$ to $J^{\prime}\approx 2$. This calibration aligns the energy gain from creating a new dimer with the energy penalty of removing two existing dimers, as one expects from local energy conservation arguments in RVB or dimer models. It injects loop resonances around defect triangles, encouraging the system toward spin-liquid-like superposition as showcased in Figure-1.
This localized increase in coupling aims to strongly induce dimerization within these specific triangles, making it energetically favorable for our ansatz to form a singlet. This targeted modification results in an average interaction strength of $J_{\text{avg}}=\approx\frac{4}{3}$ on these defect triangles. By energetically penalizing non-dimerized configurations in these frustrated regions, we compel the VQE to explore quantum states which can lead VQE closer to the true Kagome ground-state energy. Our 125-site Kagome patch, with open boundaries in both $X$ and $Y$ directions, introduces additional finite-size effects and results in a higher (less negative) ground-state energy compared to the thermodynamic limit. However, the proposed Hamiltonian engineering focuses on enabling our ansatz to efficiently discover a low-energy configuration within this finite system, thereby approximating the physics of the larger Kagome lattice.
The calibrated Hamiltonian approach leverages a hardware-efficient Variational Quantum Eigensolver (VQE) framework. Our ansatz is the Qiskit [31] Real Amplitude circuit. Figure-6(a) shows that it is 1-D chain of CNOT entanglers interleaved with parameterized single-qubit rotations, designed for compatibility with IBM’s heavy-hex superconducting architecture. This circuit can entangle nearest-neighbor qubits and is sufficiently expressive to capture local resonance physics, while remaining shallow and hardware-friendly.
Table 1: Effectiveness of Hamiltonian engineering with $J^{\prime}\approx 2$ in matching the estimated gs_energy  [^2] This data was obtained using Qiskit simulations to the exact value for progressively larger subregions of the Kagome lattice, as shown in Figure-7.
No. of Sites Subregions of Kagome Lattice of Figure-7 No. of Defect Triangles Exact gs_energy (units of J) VQE computed gs_energy (units of $J$) $J$ (units of $J$)
@@ III.2 Scalability of Hamiltonian Calibration
Table 2 summarizes example Hamiltonian engineering for various Kagome lattice subregions containing 6 to 23 sites. For each subregion size $N$, classical VQE simulations with $J^{\prime}\approx 2$ yield ground-state energies in close agreement with exact diagonalization. The corresponding geometries are depicted in Figure-7, each chosen to ensure that at least one defect triangle is present and that the ansatz qubit mapping accommodates local entanglement. In some configurations (e.g., $N=23$), two or more defect triangles are included. In each sublattice, the specific bond on the defect triangle where the Heisenberg coupling was approximately doubled is marked as $J^{\prime}$. The data highlights that the same calibrated interaction strength $J^{\prime}$ suffices across different lattice geometries without the need for re-tuning, reinforcing the scalability of the method.
@@
III.3 Physical Interpretation of $J^{\prime}\approx 2$
The smallest loop in the Kagome lattice that allows dimer flipping is the hexagonal plaquette contained within a single unit cell. We justify setting $J^{\prime}\approx 2$ by demonstrating that, at this value, the ansatz circuit enables a complete loop-flip and reproduces the exact ground-state energy. Figure-7(c) shows a 18-site Kagome unit cell whose gs_energy was computed from loop version of 6-qubit hardware-efficient ansatz circuit, which applies CNOT 5,0 to complete the loop of entangling gates. The initial dimer configuration places static singlets on edges (0,1), (2,3), and (4,5), leading to an overestimated energy of gs_energy $=-9$, while leaving three defect triangles. When the exchange coupling $J^{\prime}=2$ is applied on bonds (1,2), (3,4), and (5,0), the ansatz induces a loop-flip within the central hexagon and recovers the exact ground-state energy gs_energy $=-18$. This proof-of-concept validates that targeted Hamiltonian engineering on defect triangles can correct energy errors arising from suboptimal qubit mappings or limited ansatz expressiveness.
Table 2: Components of the gs_energy computed using local VQE  [^4] This data was obtained using Qiskit simulations. Each component corresponds to a subregion shown in Figure 8, with its respective defect triangle bond and its fine-tuned $J^{\prime}$.
Lattice subregion of the Figure-7 Bonds with $J^{\prime}$ $J$ (units of $J$) Local VQE Sub-lattice energy (units of $J$)